A park has an area of 64 square meters and a perimeter of 40 meters. Please tell me the length and width of the park.
Thanks,
Deana

Answer by checkley77(12844)

LW=64 L=64/W
2L=2W=40
2(64/W)+2W=40
128/W+2W=40
128/W=40-2W CROSS MULTIPLY.
128=W(40-2W)
128=40W-2W^2
2W^2-40W+128=0
2(W^2-20W+64)=0
2(W-16)(W-4)=0
W-16=0
W=16 ANS.
W-4=0
W=4 ANS.
L=64/16=4 ANS
L=64/4=16 ANS
PROOF:
2*16+2*4=40
31+8=40
40=40
-----------
16*4=64
64=64